OpenStudy (anonymous):

Medal+fanning What are the solutions of 6x2 − x − 2 = 0?

OpenStudy (wolf1728):

6x2 − x − 2 = 0 Use quadratic formula a=6 b=-1 c=-2 x = [1 +-sqrt(1 - 4*6*-2)]/ 2*6 x = [1 +-sqrt(49)] / 12 x = 8 / 12 = 2/3 x = 1 -7 / 12 x = -6/12 = -.5

OpenStudy (anonymous):

I think the answer is 1/2 and -2/3 but people are saying it's not that but the thing is these are the answers to this question [9.02] What are the solutions of 6x2 − x − 2 = 0? x = 1 over 2, x = −2 over 3 x = −1 over 2, x = 2 over 3 x = 1 over 6, x = −1 x = −1 over 6, x = 2

OpenStudy (wolf1728):

Actually it's -(1/2) and 2/3

OpenStudy (wolf1728):

However if the equation is changed to 6x² +x -2 = 0 Then the answers are 1/2 and -(2/3)